consecutive numbers whose product exceeds the square of the smaller by 7 ...whnat”s the answer ???

Grade:10

8 Answers

Ridhiksha hemanth mulya
8 years ago
?????
 
noogler
8 years ago
7,8
noogler
8 years ago
consecutive no. be x,x+1
given x(x+1)=x2+7..........solve it
noogler
8 years ago
approve if usefull!!!!!!!!!!!
VAISHNAVI BATHINA
8 years ago
7,8
V. Harshit -
8 years ago
7,8
x(x+1)=x^2+7
x^2+x-x^2=7
x=7
x+1=8
x and x+1are the 2 nos
Dweepayan Biswal
6 years ago
7,8
vijay kale
6 years ago
Let the numbers be x and x+1.Therefore (x+1)x=x^2+7x^2 +x=x^2+7×=7×+1=87,8 are the required number